OpenStudy (anonymous):

Pls check: use rational exponents to write single radical expression: ^4√8 * ^2√7 I have four ans to choose from: I came up with ^4√392

OpenStudy (anonymous):

Pls check: use rational exponents to write single radical expression: ^4√8 * ^2√7 I have four ans to choose from: I came up with ^4√392

OpenStudy (anonymous):

you got the correct answer here fourth root(sqrt392)

OpenStudy (anonymous):

but on you selected answers i dont see any correct one

OpenStudy (anonymous):

I sent all four through the equation editor when asked to solve the problem it gives you the four choices to choose from

OpenStudy (anonymous):

\[\sqrt[4]{392}\] is the ans I have which is one of the choices

OpenStudy (anonymous):

fourth root(sqrt392)=2.10940882 by calculator,, you can check the selected answer using calculator whichever =2.10940882 is the correct one

OpenStudy (anonymous):

^4√8 * ^2√7 =4th root√8 * 2nd root√7 =4th root√8 * 4th root(7) =4th root(7√8) =4th root(14√2) =4th root19.7989898 =2.109408824